OpenStudy (anonymous):

If 2x + y = 15 and x + 2y = -3, then 2x + 2y = A) 8 B) 12 C) 18 D) 24 E) 36

OpenStudy (eric_d):

Hello

OpenStudy (eric_d):

2x + y = 15 and x + 2y = -3, Solve this simultaneously first to find value of x and y then plug in those values into 2x + 2y =

OpenStudy (eric_d):